# Psychiatry-Bench: A Multi-Task Benchmark for LLMs in Psychiatry

### v1

@@ 4.3.2 Extended Matching Item Evaluation
In the full EMI set format, scoring follows the Partial Scoring System (PCS), which assigns partial credit based on the proportion of correctly answered subquestions in a set. In contrast, the separated subquestion format evaluates each subquestion independently using a binary scoring system.
For both formats, any instance of multiple answers being given for a single question is treated as incorrect and scored as zero.
@@ 1. Partial Scoring System (PCS) - Full EMI Set Evaluation
where $N$ is the total number of EMI sets, $S_{i}$ is Number of subquestions in the $i$-th EMI set, and $c_{ij}\in\{0,1\}$: Binary indicator for whether the $j$-th subquestion in the $i$-th set is answered correctly.
The overall EMI accuracy is the average score across all sets:
@@ 2. Individual Subquestion Evaluation - Separated EMI
Let:
  • $T$: Total number of individual subquestions across all EMI sets,
  • $a_{k}\in\{0,1\}$: Binary indicator of correctness for the $k$-th subquestion when evaluated independently.
@@
where $F_{i}$ and $S_{i}$ are indicator variables for Full format response and Separated format response on question $i$, respectively. $F=1$ if the model answers question $i$ correctly in the Full format and 0 otherwise. $S_{i}=1$ if the model answers question $i$ correctly in the Separated format and 0 otherwise. $\land$: The logical AND operator.
@@
where $C$ is the number of the correct questions in both formats, $O_{F}$ are the correct questions only in Full format, and $O_{S}$ are the correct questions only in the separated format.
@@
4.3.3 LLM as a judge
To quantitatively evaluate the output quality of model-generated answers against ground truth references, we employed an automated approach leveraging an LLM as an expert judge [54]. Specifically, Llama 3 70B was configured for this evaluation task. The core objective was to compute a similarity score, ranging from 0 to 100, representing the degree of substantive congruence between the model-generated response (referred to as the Candidate Answer) and the reference ground truth (referred to as the Reference Answer). Importantly, the evaluation was designed to focus on the core clinical content, including reasoning, correctness, and completeness of the response while disregarding differences in writing style or phrasing. 10.3
Recognizing the sensitivity of LLM evaluations to prompt design, we experimented with two distinct prompts, generated by tasking two capable LLMs, specifically GPT-4.5 and gemini-2.0-pro-exp-02-05, to create instructions for calculating the design similarity score (0–100). The prompt generation strategies differ: for GPT-4.5, we guided it toward a score-based prompt, employing a point-based system for different content components. Conversely, for gemini-2.0-pro-exp-02-05, we aimed for a more logic-driven prompt that relied on illustrative examples (few-shot learning) and emphasized understanding the overall structure and meaning of the Reference and Candidate Answers.
The prompt generated by GPT-4.5 is shown in Figure 15, and the prompt generated by gemini-2.0-pro-exp-02-05 is shown in Figure 16. After comparative testing of both prompts with the Llama 3 70B judge, we selected the gemini-2.0-pro-exp-02-05-generated prompt for our final evaluation pipeline. This decision was driven by the observation that the logic-driven, example-based prompt elicited more consistent and reliable similarity scores across a diverse range of clinical tasks and model responses. These tasks included diagnostic assessment, treatment planning, follow-up care, management strategies, clinical approaches, mental health question-answering, and multi-turn clinical reasoning. The example-based prompt aligned more closely with our focus on nuanced content evaluation, as compared to the more rigid point-based system used in the GPT-4.5 prompt.
$\text{Final Score}=\frac{1}{N}\sum_{i=1}^{N}S_{i}$ (12)
@@ 4.3.4 Classification Evaluation
Since the classification task is multi-label in nature, each instance may be associated with multiple correct labels. To accurately assess model performance in this setting, we employed two complementary evaluation metrics: Subset Accuracy and Weighted-Average F1-Score. These metrics together offer a balanced view rewarding exact correctness while also accounting for partial matches and class imbalance. [55]
@@
Where: $N$ is the total number of samples, $Y_{i}$ is the set of true labels for the $i$-th sample, $\hat{Y}_{i}$ is the set of predicted labels, $\mathbf{I}(\cdot)$ is the indicator function, and equal to 1 if $\hat{Y}_{i}=Y_{i}$, and 0 otherwise.
@@
4.4 Parsing Model Outputs
@@
Advanced reasoning-oriented models, such as DeepSeek Chat and Gemini Thinking, often diverged from strict output formats. Rather than outputting a clean label or value, they frequently embedded the answer within lengthy justifications or explanations. This behavior required parsing not just the final answer, but also understanding the structure of the response and retaining the embedded reasoning what we refer to as "printing the reasoning".
@@
5 Methodology Implementation
This section outlines the comprehensive methodology used to develop, refine, and evaluate the PsychiatryBench dataset a benchmark designed to assess the clinical reasoning capabilities of LLMs in psychiatry. We describe the end-to-end pipeline, beginning with the manual curation of high-quality psychiatric question-answer (QA) pairs from evidence-based sources, followed by rigorous filtering and formatting procedures to ensure clinical accuracy and educational value. We then present the framework used to evaluate model responses across a diverse set of task types, including both structured and open-ended formats. Finally, we detail the design and implementation of specialized tasks that test distinct dimensions of psychiatric competence, such as classification, clinical reasoning, and treatment planning. This modular and scalable methodology provides a robust foundation for benchmarking LLM performance in real-world psychiatric assessment scenarios.
5.1 PsychiatryBench Development Pipeline
Our methodology begins with systematic manual extraction of content from standard evidence-based psychiatry textbooks. Subject matter experts thoroughly review and select clinical material with high educational and diagnostic value, vibrant case studies, descriptive patient scenarios, and disorder-specific treatment information. From this material, a diverse set of question-answer (QA) pairs is constructed to represent real-world psychiatric reasoning. The QA items are crafted across eleven core task types: Diagnosis, Treatment, Mental QA, Sequential QA, Management Plan, Treatment Follow-up Clinical Approach, Classification, Multiple Choice Questions (MCQ), and Extended Matching Items (EMI). These task types were chosen for their ability to reflect lay clinical reasoning rather than rote knowledge recall.
Following construction, the QA items are passed through a rigorous filtering process designed to ensure clarity, clinical validity, and educational utility. Ambiguous, overly simple, or redundant items are removed, and the remaining questions are reviewed for content balance and diagnostic diversity. These filtered QA items form the foundation of the PsychiatryBench dataset. In this phase, questions are sorted by task type and formatted consistently, allowing for structured downstream evaluation. A comprehensive data cleaning step is performed next, standardizing answer styles, resolving typographic issues, and ensuring uniform metadata. The result is a refined, high-quality QA dataset that tests not only factual recall but also diagnostic reasoning, treatment planning, and sequential clinical decision-making in psychiatry.
5.2 Filtering, Refinement, and Data Cleaning
Following construction, the QA items are passed through a rigorous filtering process to ensure clarity, clinical validity, and educational utility. Ambiguous, overly simplistic, or redundant items are systematically removed. The remaining questions are reviewed to ensure a balance of content and a diversity of represented diagnoses. These filtered QA items form the foundation of the PsychiatryBench dataset. In this phase, questions are sorted by task type and formatted consistently to allow for structured downstream evaluation. A comprehensive data cleaning step is then performed, which involves standardizing answer styles, resolving typographic issues, and ensuring all metadata is uniform. The result is a refined, high-quality QA dataset that tests not only factual recall but also diagnostic reasoning, treatment planning, and sequential clinical decision-making in psychiatry.
5.3 Prompting Strategy
All models were evaluated using a standardized zero-shot prompting approach. While the same prompt format was applied uniformly across all models to ensure fairness in comparison, prompt design was tailored to each specific task type within PsychiatryBench. For example, diagnostic reasoning tasks employed structured clinical case prompts, whereas factual knowledge tasks like Mental QA used direct definition queries. This task-specific prompting allowed us to preserve the integrity and intent of each evaluation domain while maintaining cross-model consistency.
5.4 LLM Models Implementation
We implemented and evaluated a total of 11 LLMs, comprising both general-purpose frontier models (e.g., Gemini 2.5 Pro, Claude Sonnet 4, LLaMA 3.3, QWQ-32) and medical domain-specific models (e.g., MedGemma, OpenBioLLM). These models were tested across all eleven PsychiatryBench task types to assess their performance in real-world psychiatric reasoning tasks.
Each model was prompted using standardized engineering and evaluation templates tailored to the specific requirements of each task. Model outputs were evaluated using both automatic metrics (e.g., accuracy, F1 score) and an LLM-as-a-judge framework, which relied on a reference model to score responses based on semantic similarity and clinical relevance. This dual evaluation strategy ensured consistency, accuracy, and clinical alignment across diverse task types.
5.5 LLM Performance Evaluation Framework
To evaluate the performance of LLMs, we apply each model to the finalized PsychiatryBench dataset. Each model is prompted to generate answers across all eleven selected task types. Our evaluation framework employs a multi-faceted approach tailored to the specific nature of each task. For open-ended reasoning tasks such as Diagnosis, Treatment, and Management Plan, we utilize an LLM-as-a-judge methodology. In this dual-role capacity, a designated LLM scores the generated responses from other models based on accuracy, completeness, and clinical relevance, providing a nuanced measure of semantic alignment. For structured tasks, we apply specific quantitative metrics. For instance, the Classification task is evaluated using Subset Accuracy and F1-scores to handle its multi-label nature. At the same time, multiple-choice and Extended Matching Items are scored using accuracy and custom partial scoring systems. This hybrid evaluation framework, combining qualitative LLM-based scoring with task-specific quantitative metrics, enables a robust and scalable benchmark for assessing how well LLMs understand, reason through, and respond to complex psychiatric questions. To ensure consistency, all generated outputs were evaluated using a single evaluator model (LLaMA 3.3 70B) across all tasks. Each output and its reference were scored using a prompt-generated similarity scoring rubric. The evaluator model remained constant throughout to avoid bias introduced by evaluator variation.
5.6 Task-Specific Implementations
To capture the diverse skills required for comprehensive psychiatric reasoning, we developed five specialized task implementations within the PsychiatryBench framework. Each task is designed to probe a distinct facet of clinical expertise from high-level categorization of disorders to stepwise diagnostic reasoning, from rapid knowledge recall in board-style questions to complex pattern recognition in extended matching formats. Below, we outline the goals, data sources, construction workflows, and evaluation criteria for each task type: Classification, Clinical Approach, Multiple Choice Questions (MCQ), and Extended Matching Items (EMI). This modular design enables targeted analysis of model strengths and weaknesses across different dimensions of psychiatric decision-making.
5.6.1 Classification Task
The classification component of the PsychiatryBench was developed to evaluate the ability of LLMs to categorize psychiatric disorders based on clinical descriptions, symptom clusters, and established diagnostic systems such as the DSM-5 and ICD-10. Rather than requiring exact diagnoses, this task involved structured prompts asking models to assign each case to appropriate diagnostic categories (e.g., mood disorders, psychotic disorders) and specific disorders (e.g., bipolar I disorder, schizophrenia). Clinical case samples were manually extracted from authoritative psychiatric casebooks and annotated with preliminary disorder labels. In parallel, two reference lists were created from the DSM-5-TR: one for high-level diagnostic categories and another for specific disorders. A licensed psychiatrist validated both lists to ensure clinical accuracy and alignment with formal diagnostic criteria. A mapping process was then applied to align each sample with both the validated category-level and disorder-level labels. This dual-labeling schema enabled multi-label classification and allowed models to be evaluated on their ability to reason through hierarchical diagnostic structures. Although the classification dataset was ultimately excluded from the final model evaluation to maintain focus on more reasoning-intensive tasks, it played a critical role in shaping the design and validation of other task types, such as Diagnosis, Mental QA, and Management Plan. It also provided key insights into how LLMs interpret psychiatric terminology, differentiate between overlapping disorders, and apply structured diagnostic logic capabilities that are foundational to clinical psychiatric reasoning.
5.6.2 Clinical Approach Task
The Clinical Approach task in the PsychiatryBench is designed to evaluate a model’s ability to emulate the diagnostic thought process of a psychiatrist. Unlike other tasks that target direct question-answering, this component focuses on assessing logical reasoning, hypothesis formulation, and sequential clinical interpretation based on patient case narratives. The process begins with the creation of patient history scenarios, which are either extracted from psychiatric textbooks or crafted to simulate real-world clinical cases. These cases are then used to generate open-ended clinical reasoning questions. Question generation is performed using the Gemini 2.5 Pro model, which is prompted to formulate questions that require interpretive and multi-step reasoning. Each generated question undergoes a manual validation step. Questions that are deemed clinically relevant and educationally meaningful proceed to the next phase; otherwise, they are revised or rewritten to ensure clinical quality. Once validated, the finalized cases and questions are presented to multiple LLMs. Each model is tasked with generating structured, reasoned responses to the open-ended clinical prompts. These responses are not evaluated solely for factual correctness, but also for the quality of clinical logic, appropriateness of diagnostic pathways, and overall coherence of the suggested approach. To explore the effect of prompting methodology, two parallel evaluation tracks were conducted. In the first, questions were generated using the Gemini model. In the second, the same patient cases were used without the question-generation step, allowing for a direct comparison of LLM performance with and without prompt engineering. Each model’s output was scored across criteria of reasoning depth, clinical relevance, and judgment quality, yielding a final score. This task is essential in assessing how well LLMs can replicate the nuanced reasoning process of trained psychiatrists and highlights their potential role in clinical decision support systems.
5.6.3 Multiple Choice Question (MCQ) Task
The Multiple Choice Questions (MCQ) component was developed as part of the broader PsychiatryBench dataset to reflect the structured, high-efficiency testing style commonly used in psychiatric board examinations and educational settings. These questions typically consist of a question stem followed by four to five answer options, with only one correct choice. MCQs are widely used to assess knowledge recall, clinical recognition, and treatment matching, making them a familiar tool for both learners and educators in psychiatry.
The initial pool of MCQs was manually curated from reputable psychiatry resources, including standardized exam preparation materials, question banks, and textbook-based assessments. Many of these original items were accompanied by detailed explanations or rationales justifying why the correct answer was appropriate and why distractors were incorrect. While this information is pedagogically useful, it introduces bias when used in LLM evaluation settings, where models may memorize answer justifications or be influenced by phrasing patterns. To prevent this, a dedicated preprocessing step was implemented to remove all explanatory text from each MCQ. Ensures that only the question stem and multiple-choice options were retained, creating a clean, unbiased format suitable for model interaction and evaluation.
In addition to removing explanations, all answers were standardized. Involved converting varying answer label formats (e.g., Roman numerals, complete sentences, abbreviated codes) into a consistent A/B/C/D labeling scheme. Correct answers were annotated in a separate metadata field to facilitate scoring and model training. Furthermore, ambiguous or duplicate questions were filtered out during this process to maintain quality and reduce redundancy. Although MCQs were excluded from the final model evaluation phase in favor of tasks requiring open-ended reasoning and multi-step logic, they were retained as part of the broader dataset for auxiliary evaluation and pretraining. The MCQ component also served as a structural foundation for the more complex Extended Matching Items (EMI) task, which builds on similar principles but with greater contextual complexity.
5.6.4 Extended Matching Items (EMI) Task
The Extended Matching Items (EMI) component was developed as part of the PsychiatryBench dataset to emulate a standard format used in clinical assessments, particularly within medical and psychiatric education. EMI questions are structured to assess deeper clinical reasoning by presenting a list of possible answer options (the theme), followed by multiple clinical scenarios that require selecting the most appropriate answer from that list.
Each EMI question set typically consists of the following elements:
  • A theme (classes of antidepressants, types of psychotherapies)
  • An option list (usually 8–10 related choices)
  • Several clinical vignettes, each requiring the selection of the most suitable answer from the option list
This format is handy for testing pattern recognition, differential diagnosis, and treatment decision-making in psychiatry. For instance, an EMI set might include different patient cases presenting with depressive symptoms, and the task would be to select the most appropriate treatment or classification subtype for each case. In our methodology, EMI questions were manually created or adapted from authoritative psychiatry textbooks by expert reviewers. Ensures that each item was clinically relevant, unambiguous, and reflected real-world psychiatric decision-making. The EMI format requires the language models not only to recognize correct terminology but also to apply contextual reasoning.

### v2

@@ 4.2.2 Extended Matching Item Evaluation
In the full EMI set format, scoring follows the Partial Scoring System (PCS), which assigns partial credit based on the proportion of correctly answered subquestions in a set. In contrast, the separated subquestion format evaluates each subquestion independently using a binary scoring system. For both formats, any instance of multiple answers being given for a single question is treated as incorrect and scored as zero. the evaluation The evaluation is done as follows:
@@ 1. Partial Scoring System (PCS) - Full EMI Set Evaluation
where $S_{i}$ is the number of subquestions in the $i$-th EMI set, and $c_{ij}\in\{0,1\}$ is a binary indicator denoting whether the $j$-th subquestion in the $i$-th set is answered correctly.
The overall EMI accuracy is then calculated as the mean score across all EMI sets:
@@ 2. Individual Subquestion Evaluation - Separated EMI
where $T$: Total number of individual subquestions across all EMI sets and equals $\sum_{i}S_{i}$, and $a_{k}\in\{0,1\}$: Binary indicator of correctness for the $k$-th subquestion when evaluated independently.
@@
where $F_{i}$ and $S_{i}$ are indicator variables for Full format response and Separated format response on question $i$, respectively. $F=1$ if the model answers question $i$ correctly in the Full format and 0 otherwise. $S_{i}=1$ if the model answers question $i$ correctly in the Separated format and 0 otherwise. $\land$: The logical AND operator. And, $\mathbf{1}(\cdot)$ is the indicator function that returns 1 if the condition inside is true and 0 otherwise.
@@
where $C$ is the number of the correct questions, questions the correct questions in both formats, $O_{F}$ are the correct questions the correct questhetions only in Full format, and $O_{S}$ are the correct questions the correct questions only in the separated format.
@@ 4.2.3 Classification Evaluation
Since the classification task is multi-label in nature, each instance may be associated with multiple correct labels. To accurately assess model performance in this setting, we employed two complementary evaluation metrics: Subset Accuracy and Weighted-Average F1-Score. These metrics together offer a balanced view rewarding exact correctness while also accounting for partial matches and class imbalance [wu2020multilabelclassificationhammingloss].
@@
where $N$ is the total number of samples, $Y_{i}$ is the set of true labels for the $i$-th sample, $\hat{Y}_{i}$ is the set of predicted labels, $\mathbf{I}(\cdot)$ is the indicator function, and equal to 1 if $\hat{Y}_{i}=Y_{i}$, and 0 otherwise.
@@
4.2.4 LLM as a judge
To quantitatively evaluate the output quality of model-generated answers against ground truth references, we employed an automated approach leveraging an LLM as an expert judge [gu2024surveyllmasajudge]. Specifically, Llama 3 70B was configured for this evaluation task. The core objective was to compute a similarity score, ranging from 0 to 100, representing the degree of substantive congruence between the model-generated response (referred to as the Candidate Answer) and the reference ground truth (referred to as the Reference Answer). Importantly, the evaluation was designed to focus on the core clinical content, including reasoning, correctness, and completeness of the response while disregarding differences in writing style or phrasing Appendix 1.2.
Recognizing the sensitivity of LLM evaluations to prompt design, we systematically compared two distinct evaluation prompts generated by GPT-4.5 and Gemini 2.0 Pro. The GPT-4.5 prompt employed a rigid point-based scoring rubric with specific weights allocated to different content components (full prompt in appendix figure  20), while the Gemini-generated prompt adopted a logic-driven, few-shot learning approach emphasizing holistic understanding of clinical reasoning and semantic meaning (full prompt in appendix figure  21). After empirical testing with LLaMA 3.3 70B as the judge model, we selected the Gemini-generated prompt for our final evaluation pipeline. This decision was driven by its superior consistency and reliability across diverse psychiatric tasks, with the logic-driven, example-based approach eliciting more reliable similarity scores compared to the rigid, point-based system.
To identify the most reliable judge model, we conducted a comparative evaluation using three candidate LLMs: LearnLM (J1), LLaMA 3.3 70B (J2), and GPT-4o mini (J3). As detailed in Appendix 1.2 and Table 10, LLaMA 3.3 70B consistently assigned scores with higher discernment and sensitivity across diverse model outputs, demonstrating both consistency in ranking and meaningful differentiation of clinical quality. Consequently, LLaMA 3.3 70B was selected as our primary evaluator and remained constant throughout all experiments to eliminate evaluator-induced bias.
All generated outputs for open-ended tasks were evaluated using this single, standardized framework. Each output-reference pair was scored using the same prompt-generated similarity rubric to ensure methodological consistency across the entire benchmark. Individual similarity scores $S_{i}\in[0,100]$ obtained for each task instance i are then aggregated to compute the final evaluation score by taking the arithmetic mean across all evaluated instances.
$\text{Final Score}=\frac{1}{N}\sum_{i=1}^{N}S_{i}$ (14)
4.3 Implementation Details
We designed PsychiatryBench to ensure fair, consistent, and reproducible evaluation across models. This section outlines key implementation choices, including how prompts were standardized and how model outputs were parsed. These steps were essential for reliable cross-model comparison given the varied response behaviors of modern LLMs.
4.3.1 Prompting Strategy
All models were evaluated using a standardized zero-shot prompting approach. While the same prompt format was applied uniformly across all models to ensure fairness in comparison, prompt design was tailored to each specific task type within PsychiatryBench. For example, diagnostic reasoning tasks employed structured clinical case prompts, whereas factual knowledge tasks such as Mental QA used direct definition queries. This task-specific prompting allowed us to preserve the integrity and intent of each evaluation domain while maintaining cross-model consistency.
Across the benchmark, every task was framed as a combination of prompt prediction and prompt evaluation, ensuring that models were not only generating responses but also being assessed on their ability to interpret and follow clinical or knowledge-based instructions precisely.
We conducted two dedicated guidance sessions to refine the prompt templates used for each category. These sessions explored alternative phrasing, instruction clarity, and structural variations. After comparative analysis, we selected the prompt configuration that produced the highest performance while maintaining conceptual validity. Detailed descriptions of the guidance sessions, including the candidate prompt variations, are provided in Appendix  1.4.
For evaluation, certain tasks employed an LLM-as-a-judge framework to assess answer quality, particularly for open- ended reasoning or multi-step justification tasks. This approach ensured scalable and consistent scoring while reducing manual annotation demands. The judging rubric, reliability checks, and model parameters used in this process are detailed in Appendix  1.4.
4.3.2 Parsing Model Outputs
@@
Advanced reasoning-oriented models, such as DeepSeek Chat , Gemini Thinking and Sonnet 4.5 Thinking, often diverged from strict output formats. Rather than outputting a clean label or value, they frequently embedded the answer within lengthy justifications or explanations. This behavior required parsing not just the final answer, but also understanding the structure of the response and retaining the embedded reasoning what we refer to as "printing the reasoning".